Find the lateral and surface area of the regular square pyramid. Give exact answers. The length is 6 m and the height is 5 m. I need help.
Answer #1
A regular square pyramid has a square on the bottom. Each side of the square is the base of a triangle, (so there are four triangles).

The base is a square. Lateral area is the surface area of a 3D figure, but excluding the area of the base.

The area of each triangle is A=1/2(b)(h)
A=1/2(6)(5)=15
Multiply by four because there are four triangles to get 60 square units as your lateral area.

For the surface area, figure the area of the square base.
A=s^2
A=6^2=36
Add that to the lateral area (60+36) to get a surface area of 96 square units.
answered by: JazzMusick